A cylinder with a radius of 1.5 centimeters is half full of water. How much will the water level rise if a rock whose volume is 20 cubic centimeters is dropped into the water?
Found 2 solutions by solver91311, lwsshak3:
Answer by solver91311(24713)





Substitute your given value of and then solve what is left for

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by lwsshak3(11628)

A cylinder with a radius of 1.5 centimeters is half full of water. How much will the water level rise if a rock whose volume is 20 cubic centimeters is dropped into the water?
**
let x=rise in water
Total Volume=Area*Height of water=πr^2*height of water
Increase in volume=Area*rise in water
πr^2*x=20
x=20/πr^2=20/π*1.5^2=2.83
ans:
Level of the water will rise 2.83 cm